# Weyl metrizability of 3-dimensional projective structures and CR submanifolds

### v2

@@ Abstract.
A projective structure is Weyl metrizable if it has a representative that preserves a conformal structure. We interpret Weyl metrizability of 3-dimensional projective structures as certain 5-dimensional nondegenerate CR submanifolds in a class of 7-dimensional 2-nondegenerate CR structures. As a corollary, it follows that in dimension three Beltrami’s theorem extends to conformal structures, i.e. a flat projective structure is Weyl metrizable exclusively with respect to a flat conformal structure. In higher dimensions it is shown that conformal Beltrami theorem remains true as well.
@@ 1. Introduction
[…]
@@
To highlight some local aspects of these two questions, recall that on an $n$-dimensional manifold $M$, a torsion-fee linear connection $\nabla$ is locally defined by its Christoffel symbols, $\Gamma^{i}_{jk}=\Gamma^{i}_{kj},$ which are ${\frac{1}{2}}{n^{2}(n+1)}$ functions of $n$-variables. However, as we will see in (3.1), the projective class []delimited-[][$\nabla$][ ∇ ] is determined by the torsion-free part of $\Gamma^{i}_{jk}$’s. Thus, up to diffeomorphisms, projective structures locally depend on $a_{1}=\textstyle{\frac{1}{2}}{n^{2}(n+1)}-2n$ functions of $n$ variables. The metrizability problem can be expressed as an overdetermined first order PDE system $\Gamma^{i}_{jk}=\Gamma^{i}_{jk}(g)$ for a (pseudo-)Riemannian metric $g$ as its solution. Note that, up to diffeomorphisms, (pseudo-)Riemannian metrics on $M$ are determined by $a_{2}=\textstyle{\frac{1}{2}}{n(n+1)}-n=\textstyle{\frac{1}{2}}{n(n-1)}$ functions of $n$ variables. Thus, naively speaking, one expects to find at least $a_{1}-a_{2}=\textstyle{\frac{1}{2}}{n(n^{2}-3)}$ obstructions for the (local) metrizability of a generic projective structure on $M.$
@@
To find the local generality of Weyl structures note that from the definition above []delimited-[][$\nabla$][ ∇ ] is Weyl metrizable if there is a 1-form $\beta$ and a metric $g$ such that $\nabla g=2\beta\otimes g.$ It is immediate to check that $(\beta,g)$ and $(\beta+\mathrm{d}\lambda,e^{2\lambda}g)$ encode Weyl metrizability of $\nabla$ with respect to the same conformal structure. Thus, Weyl metrizable projective structures are encoded by a 1-form $\beta$ and a conformal structure $[g],$ i.e. their local generality, up to diffeomorphisms, is $a_{3}=n+\textstyle{\frac{1}{2}}{n(n-1)}-1$ functions of $n$ variables. As a result, naively speaking, one expects to find at least $a_{1}-a_{3}=\textstyle{\frac{1}{2}}{(n-2)(n^{2}+2n-1)}$ obstructions for the (local) Weyl metrizability of a generic projective structure on $M.$
@@
The naive count above suggests that all projective structures on surfaces are locally Weyl metrizable, which was shown in [Met14] by exploiting the twistor bundle of projective structures on surfaces, as a 2-disk bundle with a tautologically induced complex structure. The result follows from the one-to-one correspondence between (local) holomorphic sections of the twistor bundle of []delimited-[][$\nabla$][ ∇ ] and (local) Weyl structures for []delimited-[][$\nabla$][ ∇ ].
@@
In dimension three our naive count suggests the existence of local obstructions for Weyl metrizability. In this article the (local) existence of a Weyl structure for a 3-dimensional projective structure is shown to be equivalent to the (local) existence of certain 5-dimensional Levi nondegenerate CR submanifolds in a special class of 7-dimensional Levi 2-nondegenerate CR manifolds. The Levi nondegenerate CR submanifolds are obtained from the so-called twistor CR manifolds or the twistor bundle of 3-dimensional conformal structures which have been studied in the twistor programme [LeB84, Mas85]. The class of Levi 2-nondegenerate CR manifolds are the twistor bundle of 3-dimensional projective structures and have been studied recently [Gre21, KM24].
@@ 1.1. Outline of the article and main results
Throughout this article the Cartan geometric language will be employed to significantly simplify the constructions involved. We start § 2 with a brief definition of a Cartan geometry and the notion of Weyl structures for a special class of Cartan geometries referred to as parabolic geometries. Throughout the article, when dealing with Weyl structures for CR, projective, and conformal structures we will sometimes use the terms CR, projective, conformal connections, respectively. As a result, the Weyl structure that was described previously will also be called a conformal connection for some $[g].$ The rest of § 2 briefly recalls the treatment of 3-dimensional conformal structures of Riemannian and Lorentzian signature, denoted by $[g],$ as certain parabolic geometries and reviews the construction of their twistor bundle, also known as twistor CR manifolds, denoted by $\mathcal{T}^{[g]}$. The main observation in this section is Proposition 2.11 where it is shown that the holomorphic binary quartic of twistor CR manifolds, expressed in (2.9), is either zero or has a repeated root of multiplicity four. In Remark 2.12 we comment on another instance of such results in the context of conformal Fefferman metrics for 3-dimensional CR structures.
The Cartan geometric treatment of a 3-dimensional (oriented) projective structure, [],delimited-[][$\nabla$],[ ∇ ] , and the construction of its twistor bundle, denoted by $\mathcal{T}^{[\nabla]},$ is briefly reviewed in § 3. In particular, we show that the 7-dimensional twistor bundle, $\mathcal{T}^{[\nabla]},$ carries a Levi 2-nondegenerate CR structure of hypersurface type which is canonically determined by the underlying oriented 3-dimensional projective structure. See Definition 2.3 for the notion of 2-nondegeneracy.
The CR characterization of Weyl metrizability is carried out in § 4 leading to the following theorem.
@@
Given an oriented projective structure []delimited-[][$\nabla$][ ∇ ] on a 3-manifold $M,$ there is a one-to-one correspondence between conformal structures $[g]$ with respect to which []delimited-[][$\nabla$][ ∇ ] is Weyl metrizable and Levi nondegenerate CR submanifolds of hypersurface type that are transverse to the Levi kernel of $\mathcal{T}^{[\nabla]}$ whose fundamental binary quartic is either zero or has a repeated root of multiplicity four and satisfy (4.4) for $\varepsilon=1$ or $-1$ with respect to the canonical CR connection described in Proposition 4.4.
@@
A conformal structure in dimension $\geq 3$ has a projectively flat Weyl structure if and only if it is flat.
@@ 1.2. Conventions
[…]
The oriented projectivized cotangent bundle of $M$ is denoted by $\mathbb{P}_{+}T^{*}M=T^{*}M/\mathbb{R}^{+},$ i.e. the double cover of $\mathbb{P}T^{*}M.$ When dealing with Lorentzian conformal structures, $\mathbb{P}_{+}T^{*}M$ denotes the bundle of rays of future time-like 1-forms; this will be clear from the context.
Given a principal bundle $\mu\colon\mathcal{C}\to M,$ let $(\omega^{i},\omega^{i}_{j})$ be a coframe on $\mathcal{C}$ such that $\omega^{i}$’s are semi-basic with respect to $\mu\colon\mathcal{C}\to M,$ i.e. $\omega^{i}(v)=0$ for all $v\in\mathrm{Ker}\mu_{*}.$ We denote the frame associated to this coframe by $(\partial_{\omega^{i}},\partial_{\omega^{i}_{j}}).$ Moreover, given a function $f$ on $\mathcal{C},$ its coframe derivatives are defined by
@@
Except for Theorem 5.2, it is assumed that $0\leq i,j,k,l\leq 2$ and $[\varepsilon_{ij}]=\mathrm{diag}(1,\varepsilon,\varepsilon),$ where $\varepsilon$ is either $1$ or $-1$ depending on the signature of the conformal structure: It is 1 for the Riemannian case and $-1$ for the Lorentzian case. We use the summation convention over repeated indices and denote symmetrization and skew-symmetrization by round and square brackets around affected indices, e.g. $\alpha_{(ij)k}=\textstyle{\frac{1}{2}}(\alpha_{ijk}+\alpha_{jik})$ and $\alpha_{[ij]k}=\textstyle{\frac{1}{2}}(\alpha_{ijk}-\alpha_{jik}).$ Lastly, given a complex-valued function (or 1-form) $f=f_{1}+\mathrm{i}f_{2}$ its real and imaginary parts are denoted by $\Re f=f_{1}$ and $\Im f=f_{2}.$
@@ 2. Conformal structures and nondegenerate CR geometry
In this section we recall the notion of Cartan geometries and Weyl structures for the class of parabolic geometries. We review the Cartan geometric description of conformal structures in dimension three and the construction of their twistor bundle, also referred to as twistor CR manifolds. In 2.3 we review the definition of CR and para-CR structures of hypersurface type whose Levi bracket is nondegenerate or 2-nondegenerate. We finish this section with Proposition 2.11 where a necessary condition if found for a 5-dimensional nondegenerate CR structure to arise as a twistor CR manifold in terms of the root type of its holomorphic binary quartic.
@@ 2.1. Parabolic geometries and Weyl structures
Let $G$ be a Lie group and $P\subset G$ a Lie subgroup with Lie algebras $\mathfrak{g}$ and $\mathfrak{p}\subset\mathfrak{g},$ respectively. A Cartan geometry of type $(G,P)$ on a manifold $N,$ denoted as $(\mathcal{C}\to N,\psi),$ is a right principal $P$-bundle $\mathcal{C}\to N$ equipped with a Cartan connection $\psi\in\Omega^{1}(\mathcal{C},\mathfrak{g}),$ i.e. a $\mathfrak{g}$-valued 1-form on $\mathcal{C}$ satisfying
@@
    $\psi_{z}\colon T_{z}\mathcal{C}\to\mathfrak{g}$ is a linear isomorphism for all $z\in\mathcal{C}.$
@@
The 2-form $\Psi\in\Omega^{2}(\mathcal{C},\mathfrak{g})$ defined as
$\Psi(u,v)=\mathrm{d}\psi(u,v)+[\psi(X),\psi(Y)]\quad\text{for\ \ }X,Y\in\Gamma(T\mathcal{C}),$
is called the Cartan curvature and is $P$-equivariant and semi-basic with respect to the fibration $\mathcal{C}\to N.$
@@
A (local) Weyl structure for the parabolic geometry $(\mathcal{C}\to M,\psi)$ is a (local) $P_{0}$-equivariant section $s\colon\mathcal{C}_{0}\to\mathcal{C}$ of the principal $P_{+}$-bundle $\mathcal{C}\to\mathcal{C}_{0}$ where $\mathcal{C}_{0}\to M$ is the principal $P_{0}$-bundle defined as $\mathcal{C}_{0}=\mathcal{C}/P_{+}:=\mathcal{C}\times_{P_{+}}P.$
It is clear that a Weyl structure canonically defines a linear connection on the tangent bundle since at each point it gives adapted coframing defined up to the action of $P_{0}\subset\mathrm{GL}(n)$, i.e. a so-called $G$-structure. We refer the reader to [ČS09, Section 5] for more on Weyl structures.

### v3

@@ Abstract.
A projective structure is Weyl metrizable if it has a representative that preserves a conformal structure. We interpret Weyl metrizability of 3-dimensional projective structures as certain 5-dimensional nondegenerate CR submanifolds in a class of 7-dimensional 2-nondegenerate CR structures. As a corollary, it follows that in dimension three Beltrami’s theorem extends to conformal structures, i.e. a locally flat projective structure is Weyl metrizable exclusively with respect to a locally flat conformal structure. In higher dimensions it is shown that conformal Beltrami theorem remains true as well.
@@ 1. Introduction
A projective structure on a manifold $M,$ denoted as $\nabla$], is defined as the equivalence class of torsion-free linear connections whose geodesics coincide as unparametrized curves. The metrizability problem for $\nabla$], which goes back to R. Liouville [Lio89], asks whether $\nabla$] admits a representative $\nabla\in[\nabla]$ that is (locally) the Levi-Civita connection of some (pseudo-)Riemannian metric on $M.$ A slightly weaker question, referred to as the (local) Weyl metrizability problem, is whether (locally) there is a (pseudo-)Riemannian conformal structure $[g]$ on $M$ that is preserved by a representative $\nabla\in[\nabla],$ i.e. $\nabla_{X}g\in[g]$ for some, and therefore all, $g\in[g]$ along any vector field $X$ on $M.$
@@
To highlight some local aspects of these questions, we first recall from [EM08] that generic projective structures are not (locally) metrizable. It will be instructive to explain this fact by counting the local generality of projective structures and (pseudo-)Riemannian metrics. Recall that on an $n$-dimensional manifold $M$, a torsion-fee linear connection $\nabla$ is locally defined by its Christoffel symbols, $\Gamma^{i}_{jk}=\Gamma^{i}_{kj},$ which are ${\frac{1}{2}}{n^{2}(n+1)}$ functions of $n$-variables. However, as we will see in (3.1), the projective class $\nabla$] is determined by the torsion-free part of $\Gamma^{i}_{jk}$’s. Thus, up to diffeomorphisms, projective structures locally depend on $a_{1}=\textstyle{\frac{1}{2}}{n^{2}(n+1)}-2n$ functions of $n$ variables. The metrizability problem can be expressed as an overdetermined first order PDE system $\Gamma^{i}_{jk}=\Gamma^{i}_{jk}(g)$ for a (pseudo-)Riemannian metric $g$ as its solution. Note that, up to diffeomorphisms, (pseudo-)Riemannian metrics on $M$ are determined by $a_{2}=\textstyle{\frac{1}{2}}{n(n+1)}-n=\textstyle{\frac{1}{2}}{n(n-1)}$ functions of $n$ variables. Thus, naively speaking, one expects to find at least $a_{1}-a_{2}=\textstyle{\frac{1}{2}}{n(n^{2}-3)}$ obstructions for the (local) metrizability of a generic projective structure on $M,$ which explains the non-metrizability of generic projective structures.
@@
Regarding the question of Weyl metrizability, we first find the local generality of Weyl structures by noting that from the definition above $\nabla$] is Weyl metrizable if there is a connection $\nabla\in[\nabla],$ a 1-form $\beta,$ and a metric $g$ such that $\nabla g=2\beta\otimes g.$ It is immediate to check that $(\beta,g)$ and $(\beta+\mathrm{d}\lambda,e^{2\lambda}g)$ encode Weyl metrizability of $\nabla$ with respect to the same conformal structure. Thus, Weyl metrizable projective structures are encoded by a 1-form $\beta$ and a conformal structure $[g],$ i.e. their local generality, up to diffeomorphisms, is $a_{3}=n+\textstyle{\frac{1}{2}}{n(n-1)}-1$ functions of $n$ variables. As a result, naively speaking, one expects to find at least $a_{1}-a_{3}=\textstyle{\frac{1}{2}}{(n-2)(n^{2}+2n-1)}$ obstructions for the (local) Weyl metrizability of a generic projective structure on $M.$
@@
The naive count above suggests that all projective structures on surfaces are locally Weyl metrizable, which was shown in [Met14] by exploiting the twistor bundle of projective structures on surfaces, as a 2-disk bundle with a tautologically induced complex structure. The result follows from the one-to-one correspondence between (local) holomorphic sections of the twistor bundle of $\nabla$] and (local) Weyl structures for $\nabla$].
@@
In dimension three our naive count suggests the existence of local obstructions for Weyl metrizability. In this article the (local) existence of a Weyl structure for a 3-dimensional projective structure is shown to be equivalent to the (local) existence of certain 5-dimensional Levi nondegenerate CR submanifolds in a special class of 7-dimensional Levi 2-nondegenerate CR manifolds. The Levi nondegenerate CR submanifolds are obtained from the so-called twistor CR manifolds which are the twistor bundle of 3-dimensional conformal structures and have been studied in the twistor programme [LeB84, Mas85]. The class of Levi 2-nondegenerate CR manifolds are the twistor bundle of 3-dimensional projective structures and have been studied recently [Gre21, KM24].
@@ 1.1. Outline of the article and main results
Throughout this article the Cartan geometric language will be employed to significantly simplify the constructions involved. We start § 2 with a brief definition of a Cartan geometry and the notion of Weyl structures for a special class of Cartan geometries referred to as parabolic geometries. Throughout the article, when dealing with Weyl structures for CR, projective, and conformal structures we will sometimes use the terms CR, projective, conformal connections, respectively. As a result, the Weyl structure that was described previously will also be called a conformal connection for some $[g].$ The rest of § 2 briefly recalls the treatment of 3-dimensional conformal structures of Riemannian and Lorentzian signature, denoted by $[g],$ as certain parabolic geometries and reviews the construction of their twistor bundle, also known as twistor CR manifolds, denoted by $\mathcal{T}^{[g]}$. The main observation in this section is Proposition 2.11 where it is shown that the holomorphic/fundamental binary quartic of twistor CR manifolds, expressed in (2.9), is either zero or has a repeated root of multiplicity four. In Remark 2.12 we comment on another instance of such results in the context of conformal Fefferman metrics for 3-dimensional CR structures.
The Cartan geometric treatment of a 3-dimensional (oriented) projective structure, $\nabla$], and the construction of its twistor bundle, denoted by $\mathcal{T}^{[\nabla]},$ is briefly reviewed in § 3. In particular, we show that the 7-dimensional twistor bundle, $\mathcal{T}^{[\nabla]},$ carries a Levi 2-nondegenerate CR structure of hypersurface type which is canonically determined by the underlying oriented 3-dimensional projective structure. See Definition 2.3 for the notion of 2-nondegeneracy.
The CR characterization of Weyl metrizability is carried out in § 4, leading to the following theorem.
@@
Given an oriented projective structure $\nabla$] on a 3-manifold $M,$ there is a one-to-one correspondence between conformal structures $[g]$ with respect to which $\nabla$] is Weyl metrizable and 5-dimensional Levi nondegenerate CR submanifolds of hypersurface type that are transverse to the Levi kernel of the 7-dimensional twistor bundle, $\mathcal{T}^{[\nabla]},$ whose fundamental binary quartic is either zero or has a repeated root of multiplicity four and satisfy the non-integrability condition (4.4) for $\varepsilon=1$ or 1-1 with respect to the canonical CR connection described in Proposition 4.4.
@@
A conformal structure in dimension $\geq 3$ locally has a projectively flat Weyl structure if and only if it is locally conformally flat.
@@ 1.2. Conventions
In this article we will work locally over smooth manifolds. Given a distribution $\mathscr{D}\subset TM,$ its derived system is the distribution whose sheaf of sections is given by $\Gamma(\mathscr{D})+[\Gamma(\mathscr{D}),\Gamma(\mathscr{D})]$ and, by abuse of notation, is denoted as $\mathscr{D},\mathscr{D}].$ Similarly, given two distributions $\mathscr{D}_{1}$ and $\mathscr{D}_{2},$ we denote by $\mathscr{D}_{1},\mathscr{D}_{2}]$ the distribution whose sheaf of sections is $\Gamma(\mathscr{D}_{1})+\Gamma(\mathscr{D}_{2})+[\Gamma(\mathscr{D}_{1}),\Gamma(\mathscr{D}_{1})]+[\Gamma(\mathscr{D}_{1}),\Gamma(\mathscr{D}_{2})]+[\Gamma(\mathscr{D}_{2}),\Gamma(\mathscr{D}_{2})].$ We denote by $\Omega^{k}(M)$ and $\Omega^{k}(M,\mathfrak{g})$ the sheaf of sections of $\bigwedge\nolimits^{k}T^{*}M$ and $\bigwedge\nolimits^{k}T^{*}M\otimes\mathfrak{g},$ respectively. The symmetric product of 1-forms is denoted by their multiplication, e.g. for two 1-forms $\alpha$ and $\beta$ we define $\alpha\beta=\textstyle{\frac{1}{2}}(\alpha\otimes\beta+\beta\otimes\alpha)$ and $\alpha^{k}$ denotes the $k$th symmetric power of $\alpha.$ The span of vector fields $v_{1},\dots,v_{k}\in\Gamma(TM)$ is denoted by $\langle v_{1},\dots,v_{k}\rangle$ and the algebraic ideal generated by 1-forms $\omega^{0},\dots,\omega^{n}\in\Omega^{1}(M)$ is denoted as $\{\omega^{0},\dots,\omega^{n}\}.$
When dealing with a Lorentzian conformal structures on $M$, the bundle of rays of future time-like 1-forms is denoted by $\mathcal{R}^{\uparrow}T^{*}M$. In the case of conformal structures of Riemannian signature, $\mathcal{R}^{\uparrow}T^{*}M$ coincides with the oriented projectivized cotangent bundle $\mathbb{P}_{+}T^{*}M:=T^{*}M/\penalty 50\mathbb{R}^{+},$ which is the double cover of $\mathbb{P}T^{*}M,$
Given a principal bundle $\mu\colon\mathcal{C}\to M,$ let $(\omega^{i},\omega^{i}_{j})$ be a coframe on $\mathcal{C}$ such that $\omega^{i}$’s are semi-basic with respect to $\mu\colon\mathcal{C}\to M,$ i.e. $\omega^{i}(v)=0$ for all $v\in\mathrm{Ker}\mu_{*}.$ We denote the frame dual to this coframe by $(\partial_{\omega^{i}},\partial_{\omega^{i}_{j}}).$ Moreover, given a function $f$ on $\mathcal{C},$ its coframe derivatives are defined by
@@
Except for Theorem 5.2, it is assumed that $0\leq i,j,k,l\leq 2$ and $\varepsilon_{ij}]=\mathrm{diag}(1,\varepsilon,\varepsilon),$ where $\varepsilon$ is either $1$ or 1-1 depending on the signature of the conformal structure: It is 1 for the Riemannian case and 1-1 for the Lorentzian case. We use the summation convention over repeated indices and denote symmetrization and skew-symmetrization by round and square brackets around affected indices, respectively, e.g. $\alpha_{(ij)k}=\textstyle{\frac{1}{2}}(\alpha_{ijk}+\alpha_{jik})$ and $\alpha_{[ij]k}=\textstyle{\frac{1}{2}}(\alpha_{ijk}-\alpha_{jik}).$ Lastly, given a complex-valued function (or 1-form) $f=f_{1}+\mathrm{i}f_{2}$ its real and imaginary parts are denoted by $\Re f=f_{1}$ and $\Im f=f_{2}.$
@@ 2. Conformal structures and nondegenerate CR geometry
In this section we recall the notion of Cartan geometries and Weyl structures for the class of parabolic geometries. We review the Cartan geometric description of conformal structures in dimension three and the construction of their twistor bundle, also referred to as twistor CR manifolds. In § 2.3 we review the definition of CR and para-CR structures of hypersurface type whose Levi bracket is nondegenerate or 2-nondegenerate. We finish this section with Proposition 2.11 where a necessary condition is found for a 5-dimensional nondegenerate CR structure to arise as a twistor CR manifold in terms of the root type of its holomorphic/fundamental binary quartic.
@@ 2.1. Parabolic geometries and Weyl structures
Let $G$ be a Lie group and $P\subset G$ a Lie subgroup with Lie algebras $\mathfrak{g}$ and $\mathfrak{p}\subset\mathfrak{g},$ respectively. A Cartan geometry of type $(G,P)$ on a manifold $N,$ denoted as $(\mathcal{G}\to N,\psi),$ is a right principal $P$-bundle $\mathcal{G}\to N$ equipped with a Cartan connection $\psi\in\Omega^{1}(\mathcal{G},\mathfrak{g}),$ i.e. a $\mathfrak{g}$-valued 1-form on $\mathcal{G}$ satisfying
@@
    $\psi_{z}\colon T_{z}\mathcal{G}\to\mathfrak{g}$ is a linear isomorphism for all $z\in\mathcal{G}.$
@@
The 2-form $\Psi\in\Omega^{2}(\mathcal{G},\mathfrak{g})$ defined as
$\Psi(X,Y)=\mathrm{d}\psi(X,Y)+[\psi(X),\psi(Y)]\quad\text{for\ \ }X,Y\in\Gamma(T\mathcal{G}),$
is called the Cartan curvature and is $P$-equivariant and semi-basic with respect to the fibration $\mathcal{G}\to N.$
@@
A (local) Weyl structure for the parabolic geometry $(\mathcal{G}\to M,\psi)$ is a (local) $P_{0}$-equivariant section $s\colon\mathcal{G}_{0}\to\mathcal{G}$ of the principal $P_{+}$-bundle $\mathcal{G}\to\mathcal{G}_{0}$ where $\mathcal{G}_{0}\to M$ is the principal $P_{0}$-bundle defined as $\mathcal{G}_{0}=\mathcal{G}/\penalty 50P_{+}:=\mathcal{G}\times_{P_{+}}P.$
Given a Weyl structure $s\colon\mathcal{G}_{0}\to\mathcal{G},$ let $\psi_{0}$ be the $\mathfrak{g}_{0}$-valued part of $s^{*}\psi.$ It is clear that, given a Weyl structure, $\psi_{0}$ defines a linear connection on the tangent bundle for the underlying $G$-structure where $G=P_{0}\subset\mathrm{GL}(n)$, which is referred to as a Weyl connection. We refer the reader to [ČS09, Section 5] for more on Weyl structures.